# Stateful Large Language Model Serving with Pensieve

### v1

@@ 5. Implementation
We have implemented our prototype serving system Pensieve with $\sim$5K lines of C++/CUDA code. Pensieve relies on PyTorch C++ API front-end to perform GPU Tensor memory management and to execute Large Language Models. Based on PyTorch’s fused memory efficient attention, we develop our own multi-query fused attention kernel using NVIDIA Cutlass library to support performing attention with KV-tokens that reside in non-contiguous memory.
5.1. Optimization: Prioritize Data Transfer to GPU over eviction to CPU
Although PCIe sends data full-duplex bidirectionally, in practice, we found that when CPU-to-GPU transfer is overlapped with GPU-to-CPU transfer, there is a significant throughput drop (18-20%) in both directions. Similar issues have been reported [^2] https://forums.developer.nvidia.com/t/data-transfers-are-slower-when-overlapped-than-when-running-sequentially/187542. Since Pensieve performs KV-token backup ahead of time, there is no urgency to finish the transfer right away. In order to prevent evicting from slowing down swapping in past KV-token, we set up a blocking mechanism. Whenever worker plans to issue GPU-to-CPU copy, it first checks if there is any swap-in task in the system, and waits until swap-in task is done. Although this conservative approach does not fully utilize duplexed PCIe bandwidth, due to the existence of generation phase, we find that this optimization performs well and we never run into the situation that the backup can’t catch up.
@@ 6. Evaluation
In this section, we show the evaluation results of Pensieve.
Model size 6.7B 13B
# layer 32 40
# hidden 4096 5120
# head 32 40
Layer MFLOPs Linear 402.65 629.15
Attention 33.55 41.94
Total 426.20 671.09
Table 1. Evaluated GPT-3 model configurations, and approximate amount of computation (FLOPs) in each Transformer layer normalized by the number of input sequence tokens.
ShareGPT LMSYS
# conversations 15553 79166
Average # turns 7.58 7.73
Average input length 31.26 22.49
Average output length 180.44 100.67
Table 2. Dataset staticstics
(a) GPT3-6.7B, ShareGPT
(b) GPT3-13B, ShareGPT
(c) GPT3-6.7B, LMSYS
(d) GPT3-13B, LMSYS
Figure 8. 90 percentile normalized latency vs throughput graph for GPT-3 model on ShareGPT and LMSYS datasets.
@@ 6.1. Experimental Setup
System Environment
We evaluate Pensieve on a Microsoft Azure NC24ads_A100_v4 instance, which is equipped with 24 non-multithreaded AMD EPYC Milan processor cores, 220 GB system memory, and one NVIDIA A100 PCIe GPU with 80GB memory. We configure each system to allocate 40GB GPU memory to KV cache for fair comparison. Pensieve and all baselines use CUDA version 11.8 and PyTorch 2.0.
Model
We use the popular GPT-3 (gpt3, ) as our benchmark model and evaluate two configurations listed in the GPT-3 paper: 6.7 billion parameters and 13 billion parameters. Detailed model hyper-parameters can be found in Table 1. For all experiments, 16-bit half-precision float format is used for both model parameters and intermediate hidden representations. Model parameters are randomly initialized.
@@
We evaluate Pensieve using two datasets collected from AI-powered ChatBot: ShareGPT (sharegpt, ), a dataset containing user-shared ChatGPT conversations [^3] We used this version: https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, which performed additional filtering like removing HTML tags and non-English conversation., and LMSYS-Chat-1M (lmsys, ). Since we target multi-turn conversation scenario, we only keep conversations with at least 5 turns. We also limit the maximum context size to 2048 (larger context size studied in §6.3). Statistics about datasets are listed in Table 2. We extract the input and output lengths of each conversation turn from these datasets to construct requests. To ensure the same computation amount, we ask each system to ignore any early stop criteria, e.g. end-of-sentence (EOS) token, and keep generating until the ground-truth output length is reached.
@@
Since the dataset does not provide timestamps for when each user prompt is received, we sample arrival time for each request from Poisson distribution under different request rates. We maintain the causal dependency for requests belonging to the same conversation: a new user prompt is only sent to the system after earlier ones from the conversation have already received their response. In addition, in real-life chatting scenarios, it takes time for users to reply. We assume each generated output token adds 100 milliseconds to such user response time. For example, if 200 tokens are generated for a conversation, then within the next 20 seconds, no new request will come from this conversation. Such response time is set to have a minimum value of 5 seconds.
Baseline
We evaluate our system against the state-of-the-art vLLM serving system (vllm, ) (version v0.2.0), which uses a stateless serving API. For each request, the user needs to concatenate chat history with the new user question as input prompt. Since vLLM is a PyTorch-based execution engine with PagedAttention kernel integrated, and has no perpetually serving loop to drive the engine to continue execution, we implement such a driver that adds newly arrived requests into vLLM job queue and invokes the engine execution until all requests are fully processed.
We also implement a variant of Pensieve called Pensieve (GPU cache) that does not use CPU memory and only saves past KV-tokens in spare GPU memory spaces to reuse them in the future.
@@
Pensieve is designed to optimize both peak throughput and latency of serving LLM in conversational scenarios. Following prior work (orca, ; vllm, ), we measure the 90 percentile normalized latency and the achieved request serving throughput. Normalized latency is calculated as the end-to-end request serving latency divided by the number of output tokens.
6.2. End-to-end Pensieve Performance
Figure 8 is the 90 percentile latency (normalized by request output length) vs throughput graph of the end-to-end serving performance of Pensieve and baselines. When compared against vLLM, for GPT3-6.7B and GPT3-13B, Pensieve is able to achieve a throughput gain of 1.51x and 1.65x on ShareGPT dataset, and 1.84x and 1.95x on LMSYS dataset. At the point when vLLM baseline is close to its peak throughput, Pensieve is able to reduce 90 percentile normalized latency by about 60-75%.
The throughput gain on LMSYS dataset is in general larger than ShareGPT dataset. The reason is that requests in LMSYS dataset have on average smaller output length (see Table 2), and as a result, the amount of time spent on generation phase is shorter. Since Pensieve focuses on optimizing prompt prefill phase by avoiding recomputing past KV-tokens, a dataset with a relatively longer prefill phase favors our solution.
Experiments using GPT3-13B achieve about 12% higher throughput gain on both datasets when compared against GPT-6.7B. This is due to the fact that as the model dimension increases, the computation amount increases faster than the context cache size that we need to transfer from CPU. The bottom part in Table 1 lists the computation amount by major operations in each Transformer Layer. Linear layer operations  [^4] Despite having more FLOPs in absolute number, linear layer operators can make use of highly efficient NVIDIA Tensor Cores. grows quadratically with respect to the growth of hidden size. When the model hidden size is increased from 4096 to 5120, which amounts to 25% more data to copy, linear operation’s computation amount increases by around 50%. Therefore, recomputing full context from scratch becomes more costly. With the trend of increasingly larger models, we expect to see Pensieve becoming even more beneficial.
Pensieve (GPU cache), the yellow line in Figure 8, demonstrates the performance of using GPU cache only without evicting to CPU. This approach utilizes unoccupied GPU space to keep past KV-tokens of conversations for as long as possible, and only discard them when GPU KV cache is out of memory. When the system is serving under a low request rate, GPU-side caching only may also deliver a similar serving latency of Pensieve. However, as workload rates go up, more conversations are being served at the same time, and GPU memory space exhausts quickly. Eventually, at a higher rate, this approach behaves similarly to recomputing past KV-tokens from scratch.
6.3. Larger maximum context size
(a) ShareGPT
(b) LMSYS
Figure 9. 90 percentile normalized latency vs throughput graph for GPT3-13B model with maximun context size 16384
In scenarios where the conversation lasts lots of turns, or the user feeds a long article for the AI agent to read and answer questions, context size may exceed the commonly used 2048 limit. Figure 9 shows how longer context affects the performance. From the figure, with maximum context size 16384, the throughput gain Pensieve can achieve over the baselines decreases. On one hand, longer conversations need more memory to store. On the other hand, longer conversations tend to have more turns, last longer, and have to be stored longer, making it more likely for the system to experience KV-token drop and recomputation.
6.4. Efficiency of multi-query attention kernel
@@
Figure 10. Execution time of Pensieve’s multi-query attention kernel over non-contiguous context memory with different context size. Measured with batch size 32 and query size 8.
@@
Figure 10 shows the advantage of Pensieve’s multi-query attention kernel when used to prefill KV-tokens of prompts from new requests with their past KV-tokens. In this microbenchmark, a batch of 32 requests perform attention between a prompt of 8 query tokens and different numbers of past KV-tokens stored in non-contiguous GPU memory. As described in §3.1, existing attention kernels are not directly applicable. We compare against two straw-man implementations: 1) allocate additional contiguous GPU memory space to copy out past KV-tokens and then apply a fused attention kernel (orange bar), and 2) apply multiple rounds of vLLM’s single-query PagedAttention to process one token from the prompt at a time (green bar). We also present the performance of the ideal situation which assumes that past KV-tokens are already in contiguous memory space (blue bar). From Figure 10, both straw-man solutions add significant overhead compared to the ideal performance. Copy out requires extra memory space and gathers past KV-tokens from non-contiguous space. Such cost is proportional to the number of past KV-tokens. Applying multiple rounds of PagedAttention, on the other hand, gives up parallelization opportunities on prompt tokens, leading to linear execution time with respect to the number of tokens in the prompt. Pensieve’s kernel shows slightly better performance than the ideal baseline because we offload auxiliary data computing (like calculating the cumulative sum for the sequence length of a batch) to CPU. Since each transformer layer in the model shares the same caching plan, these auxiliary data can be reused by all layers.
7. Discussion
Distributed Execution
For large LLM model that needs to be sharded across multiple GPUs, Pensieve follows Model Parallelism (also called Tensor Parallelism) pattern of Transformer layers proposed by NVIDIA Megatron-LM (megatron, ). With model parallelism, each GPU worker executes exactly the same computation and therefore, the same context cache arrangement can be shared by all workers.
Distributed version Pensieve is not fully vetted yet. We will provide concrete evaluation results for multi-GPU and larger models in next version of this paper.
@@ 8. Related Works
8.1. Systems Tailored for Serving Language Models
Many systems have been designed to serve language models with high performance. For example, systems like FasterTransformer (fastertransformer, ), LightSeq (lightseq, ), DeepSpeed (deepspeedinference, ), etc., adopt incremental decoding to avoid duplicated computation, provide better fused kernels and quantization support, and improve the efficiency of decoding algorithms like beam search and various sampling algorithms.
A key challenge presented by language model serving is that requests may have different input prompt sizes, and the number of output tokens to generate is unknown beforehand until stopping criteria is met. Most systems handle this variable length issue by padding each request in the batch to the same lengths. In order to make more efficient batching decisions, BatchMaker (batchmaker, ) proposed to use finer batching granularity and perform scheduling at the granularity of the token. ORCA (orca, ) further extended this idea to Transformer-based models by adding support to attention operations.
Among all systems serving language models, vLLM (vllm, ) is the most closely related one to this work, as it focuses on LLM KV cache memory management. vLLM addresses memory fragmentation caused by variable request prompt size and unknown generation size by adapting virtual memory paging solution from the operating system. It divides memory into pages, each holding a fixed number of tokens. Instead of allocating memory to maximum context size, vLLM allocates physical GPU memory on demand and enables memory sharing among requests.
8.2. Techniques Targeting Limited GPU Memory
Many techniques address GPU’s limited memory space, including swapping, recomputation, unified memory, etc.
Swapping / offloading
SwapAdvisor (swapadvisor, ) performs GPU-CPU swapping for DNN training. It analyzes dataflow graphs and decides operator execution order, memory allocation, and the swapping plan. Zero-Offload (zerooffload, ) offloads optimizer states and gradients to CPU for training LLM with a single GPU. DeepSpeed-ZeroInference (deepspeedinference, ) and FlexGen (flexgen, ) further adapt the offloading techniques to serve LLM with a weak GPU. DeepSpeed-ZeroInference offloads entire model weights to CPU or NVMe memory. FlexGen offloads currently unused model weights, activation or KV cache to CPU memory or disk. But as a result of frequent data movement and high disk latency, these two systems require a large batch size so that computation amount is large enough to hide the offloading latency. Therefore, they mainly target latency-insensitive applications.
Recomputation
In training, model activation recomputation (recompute, ) is a commonly used technique to reduce memory footprint. A subset of activations with relatively low computation cost is selected to be dropped after being used in model forward pass and recomputed on demand in backward pass. Checkmate (checkmate, ) further formalizes the trade-off between time and memory usage as tensor rematerialization optimization problem and finds the optimal schedule using MILP solver. (reducing, ) exploits the Transformer layer structure to only selectively recompute attention module which has larger memory consumption but low number of floating-point operations to improve the cost efficiency of trading recomputation time for memory space. But in inference setting, we only preserve the Key Value embedding of tokens for incremental decoding and all other activations are discarded. KV embeddings are very costly to recompute, and therefore, we opt to swap them to CPU.
Unified Memory and Direct Host Access
Moving LLM KV cache between CPU and GPU memory can also be achieved in a transparent manner through Unified Memory (unifiedmemory, ), which, similar to CPU virtual memory, allows users to oversubscribe GPU memory and automatically triggers memory page migration between CPU and GPU using a page fault mechanism. However, CUDA Unified Memory may cause longer latency due to the overhead of page migration. DeepUM (deepum, ) addresses this issue by developing better prefetching techniques. HUVM (huvm, ) further provides unified multi-GPU virtual memory and allows transparently storing data on neighboring GPUs to utilize the high-bandwidth interconnect among GPUs to reduce access latency caused by host memory access via slower PCIe.
Another line of works (direct, ; deepplan, ) utilizes Direct Host Access feature of modern GPU, which allows GPU kernel to directly read from host memory. The main difference compared to CPU to GPU memory copy (explicit CUDAMemcpy or implicit Unified Memory) is that the data transfer to GPU is immediately consumed by the GPU thread requesting it. This brings better overlapping of data transfer and computation because part of threads in GPU kernel can start to execute earlier without waiting for the entire transfer to finish as long as the data these threads requested are ready in GPU.
Neither Unified Memory nor Direct Host Access is the most efficient choice to solve our KV cache management problem, because their memory transfer is only triggered when data is accessed by a GPU kernel, which unnecessarily gives up the opportunity to fetch ahead of time. Hence, we decided to manage data movement explicitly. Furthermore, for Direct Host Access, since the copied data is not explicitly stored in GPU memory, it has to be loaded from CPU again if it is repeatedly used. In our workload, the context cache is not only used to prefill the prompt tokens but also in every generation step, which makes Direct Host Access significantly less efficient.
8.3. Techniques that speed up LLM inference
Kernel Fusion Deep Neural Networks often have a lot of memory-bound operations, e.g. element-wise operators, that read a large amount of data while performing little computation. A common technique to improve GPU efficiency is kernel fusion (xla, ; tvm, ), which fuses multiple kernels as one single kernel to avoid storing intermediate data into GPU global memory and later loading them back. Attention in LLM can also benefit from kernel fusion. Memory efficient attention (memeff, ) and Flash Attention (flash, ) fuse all attention-related operations into one kernel to avoid materializing the attention score tensor. The attention kernel we implemented in Pensieve is a generalization of above mentioned fused kernels, with support to efficiently attend new tokens with past KV-tokens stored in non-contiguous memory. Recently, Flash-decoding (flashdecode, ) further parallelizes the processing of KV-tokens for single query token attention, which can significantly speed up the generation phase. With prefill phase taking up a larger portion of total computation, Pensieve’s optimization for prefill phase may become even more beneficial.
Quantization Quantization techniques use lower precision floating point or integer representation to speed up execution and reduce memory footprint (nagel2021white, ; gholami2022survey, ). 16-bit half precision is widely adopted in LLM inference and mixed precision training (mixed, ). And recently, several works (dettmers2023case, ; qlora, ; flexgen, ) have shown that quantizing to 4-bit can still maintain good LLM inference quality. In this work, we focus only on 16-bit half-precision to deliver unchanged inference results. But in general, quantization-based methods are also applicable to our system.
Speculative Decoding Recently, speculative decoding (speculative:chen, ; speculative:leviathan, ) is proposed to speed up the generation phase in LLM inference. Instead of running a costly LLM model to sequentially produce 1 output token in each model iteration, speculative techniques use a smaller and faster model to generate a draft of a series of tokens, which get evaluated by a larger and more powerful LLM model later with more efficient parallel execution. The adoption of speculative decoding significantly reduces the execution time of generation phase, which makes Pensieve’s optimization of prefill phase in multi-turn conversation serving more important.
8.4. Other General DNN Serving Systems
@@
There are many general-purpose systems, like TensorFlow Serving (tfserving, ), Clipper (clipper, ), NVIDIA Triton Inference Server (triton, ), Nexus (nexus, ), InferLine (inferline, ) that serve as scheduling component of the serving system. While being mostly model-agnostic and execution backend agnostic, these systems apply general system techniques like batching, caching, and software pipelining to serve Deep Learning models, and also properly provision compute resources to improve overall cluster efficiency. On the other end of the spectrum, several works target model-less serving. They aim to provide inference as a service and automatically select models to meet accuracy and latency requirements. For example, INFaaS (infaas, ) and Tabi (tabi, ) serve requests with small model and only re-route to larger models when the output confidence score is low.

### v2

@@ 5. Implementation
We have implemented our prototype serving system Pensieve with $\sim$7K lines of C++/CUDA code. Pensieve manages KV cache and auxiliary data structure needed by multi-token attention on the GPU, but relies on PyTorch (v0.2.0, CUDA 11.8) C++ front-end APIs to execute GPU operators in Large Language Models. Based on PyTorch’s fused memory efficient attention, we develop our own fused multi-token attention kernel using NVIDIA Cutlass library to support performing attention with KV-tokens that reside in non-contiguous memory.
Optimization: Prioritize data retrieval over eviction
Although PCIe allows full-duplex bidirectional data transfer, in practice, we found that when CPU-to-GPU data transfer is done concurrently with GPU-to-CPU data transfer, there is a significant throughput drop (18-20%) in both directions. Similar issues have been reported [^2] https://forums.developer.nvidia.com/t/data-transfers-are-slower-when-overlapped-than-when-running-sequentially/187542. Since Pensieve performs KV-token swap-out ahead of time, there is no urgency to finish the transfer right away. To prevent eviction from slowing down the swapping in of past KV-token, we set up a waiting mechanism. In particular, if a worker has any ongoing swap-in task, it waits to perform GPU-to-CPU copy until the swap-in task is done. Although this conservative approach does not fully utilize the duplex PCIe bandwidth, we find that this optimization performs well and we never run into the situation that the GPU-to-CPU copying can’t catch up.
@@ 6.1. Experimental Setup
System Environment.
We evaluate Pensieve on Azure NC A100 v4 series, which are equipped with up to 4 A100-80GB GPUs, a 24-core AMD EPYC 7003 processor per GPU, and 220 GB CPU memory. For each system evaluated, we configure it to use 40GB GPU memory for KV cache for a fair comparison.
Models
Model OPT-13B OPT-66B Llama 2-13B Llama 2-70B
# layer 40 64 40 80
# hidden 5120 9216 5120 8192
# head 40 72 40 64
# KV head 40 72 $10^{\ast}$ 8
Head size 128 128 128 128
# GPU 1 4 1 4
Table 1. Hyper-paramaters for OPT and Llama 2 models.
We use two open-source models: OPT (Zhang et al., 2022) and Llama 2 (Touvron et al., 2023b). OPT has an almost identical model architecture to GPT-3 (Brown et al., 2020) while Llama 2 is a more recent model that employs more advanced model features like rotary embedding, RMS Layernorm (Zhang and Sennrich, 2019), SiLU, etc. Notably, Llama 2 follows the trend of adopting Grouped-Query Attention (GQA) (Ainslie et al., 2023) which divides query heads into groups so that only one KV head is used within each group. GQA significantly reduces the memory consumption of KV-tokens, allowing Pensieve to store more past KV-tokens.
We evaluate two different sizes for each model: a small one on a single GPU, and a large one partitioned onto 4 GPUs using Tensor Parallelization as done in Megatron-LM (Shoeybi et al., 2019). Detailed model hyper-parameters can be found in Table 1. By default, Llama 2 only uses GQA for models with over 70 billion parameters. To demonstrate Pensieve’s effectiveness when used with GQA, we changed the number of KV heads of Llama 2-13B from 40 to 10. In all experiments, the 16-bit half-precision float format is used for both model parameters and intermediate hidden representations.
@@
We evaluate Pensieve on two multi-turn conversation datasets: ShareGPT and UltraChat. ShareGPT (ShareGPT, 2023) is a real-world dataset containing user-shared ChatGPT conversations. UltraChat (Ding et al., 2023) is a recent large-scale synthetic dataset for multi-turn dialogue: it uses separate LLMs to simulate the interaction between a user and the chatbot assistant. Table 2 shows the statistics of both datasets. In our experiments, we limited the maximum context size to 16384 tokens and dropped 0.57% of the conversations in ShareGPT dataset that exceed this limit.
ShareGPT UltraChat
# conversations 48,159 1,468,352
Mean # of turns 5.56 3.86
Mean request input length 37.77 51.78
Mean request output length 204.58 257.81
Table 2. Dataset statistics
@@
Since the datasets do not provide timestamps for each user request, we simulate a request’s arrival time by sampling from a Poisson distribution under different request rates. We maintain the causal dependency for requests belonging to the same conversation: a new user prompt is only sent to the system after the response to the conversation’s previous request has been received. Additionally, we also simulate user think time, aka the time taken for users to generate the next conversation turn, by sampling from an exponential distribution with varying mean.
Baselines
We evaluate Pensieve against two state-of-the-art serving systems: vLLM (Kwon et al., 2023) (v0.2.0) and TensorRT-LLM (NVIDIA, 2023) (v0.9.0). Both of them use a stateless serving API. For each request, the new user prompt is appended to the history and then processed as an input request. vLLM uses PyTorch as its execution backend. By contrast, TensorRT-LLM compiles and optimizes the model using graph rewriting optimizations such as operator fusion, and executes the optimized model using the TensorRT Runtime. Since vLLM is an execution engine without a serving loop that drives the engine to process incoming requests, we implement such a driver that adds newly arrived requests into vLLM job queue and invokes the engine execution until all requests are fully processed.
We also experiment with a variant of Pensieve called Pensieve (GPU cache) that simply drops evicted tokens from the GPU instead of swapping them out to the CPU. This variant is used to examine the effectiveness of CPU caching in Pensieve.
(a) OPT-13B, ShareGPT
(b) Llama 2-13B, ShareGPT
(c) OPT-13B, UltraChat
(d) Llama 2-13B, UltraChat
Figure 10. LLM serving performance on 1 GPU.
@@
Pensieve is designed to optimize both peak throughput and latency of serving LLM in conversational scenarios. Following prior work (Yu et al., 2022; Kwon et al., 2023), we measure the achieved serving throughput and 90-percentile normalized latency, which is calculated as the end-to-end request serving latency divided by the number of output tokens.
6.2. End-to-End Serving Performance
Figure 10 shows the normalized latency vs throughput for OPT-13B and Llama 2-13B served with a single A100-80GB GPU. The normalized latency is calculated as the mean of each request’s end-to-end latency divided by its output length, as done in (Yu et al., 2022; Kwon et al., 2023). On the ShareGPT dataset shown in Figure 10(a) and (b), Pensieve achieves 33% more throughput over vLLM and 19% more over TensorRT-LLM for serving OPT-13B, and  57% more throughput over both vLLM and TensorRT-LLM for serving Llama 2-13B. On the UltraChat dataset shown in Figure 10(c) and (d), Pensieve achieves 17% and 13% more throughput over vLLM and TensorRT-LLM respectively for serving OPT-13B, 58% and 47% more throughput for serving Llama 2-13B. Pensieve has more throughput gains on ShareGPT than UltraChat because the real world dataset ShareGPT has more conversation turns (Table 2), which makes saving past KV-tokens more beneficial.
Figure 10 shows that TensorRT-LLM outperforms vLLM. This is expected because TensorRT-LLM compiles and optimizes the model before execution. However, despite the fact that Pensieve also executes the model directly using Torch like vLLM, it outperforms TensorRT-LLM by avoiding recomputing past KV-tokens for continuing conversations.
Pensieve’s performance gain is more significant for Llama 2-13B than for OPT-13B because our version of Llama 2-13B uses Grouped-Query Attention with group size 4 (i.e. every four query heads share a single key-value head). Consequently, the amount of memory required to store past KV-tokens is reduced by 4x, thus allowing Pensieve to store more past KV-tokens and better avoid recomputation.
When Pensieve is used without CPU cache, i.e. Pensieve (GPU cache), it may still reduces latency because baseline systems always recompute past KV-tokens for requests from a returning conversation, and thus, the number of tokens processed in the prefill phase is on average larger. But under relatively higher request rates, GPU cache is quickly exhausted, and Pensieve (GPU cache) also resorts to recomputing past KV-tokens from scratch.
6.3. Multi-GPU Serving Performance
(a) OPT-66B
(b) Llama 2-70B
Figure 11. LLM serving performance on 4 GPUs for ShareGPT.
Figure 11 shows the performance of Pensieve for larger models, OPT-66B and Llama 2-70B, when run on four GPUs using the ShareGPT dataset. Larger models amplify Pensieve’s advantage over the baselines because the amount of computation grows faster than the memory usage of KV-tokens. For example, from OPT-13B to OPT-66B, the model parameter size and computation amount increase by more than 4x, while the hidden size only increases 1.8x from 5120 to 9216 (Table 1). Since the number of GPUs and CPU memory are usually scaled linearly with the model size, Pensieve can store more past KV-tokens in its KV cache.
Pensieve achieves 1.5x the throughput of baseline systems for OPT-66B and 1.6x for Llama 2-70B. The improvement is more significant on Llama 2-70B because it uses Grouped-Query Attention with group size 8, which reduces the memory requirement for past KV-tokens by 8x. This much-reduced memory consumption in KV-tokens also significantly benefits the throughput of Pensieve (GPU cache), as shown in Figure 11(b).
6.4. Performance of Multi-token attention Kernel
@@
Figure 12. Execution time of Pensieve’s multi-token attention kernel over non-contiguous context memory with different context size. Measured with batch size 32 and query size 8.
@@
Figure 12 shows the performance of Pensieve’s multi-token attention kernel compared to alternative implementations. In this microbenchmark, we measure the latency of the attention operator for a batch of 32 requests each with a prompt of 8 query tokens and different numbers of past KV-tokens stored in non-contiguous GPU memory. As described in §3.1, existing attention kernels are not directly applicable. We compare against two straw-man implementations: (1) “CopyOut+Attention” allocates additional contiguous GPU memory to copy past KV-tokens into and then invokes an existing fused attention kernel (orange bar), and 2) “Multi-round PagedAttention” invokes multiple rounds of vLLM’s single-query PagedAttention to process one token from the prompt at a time (green bar). We also show the performance of the ideal situation which assumes that past KV-tokens are stored in contiguous memory space (blue bar). Figure 12 shows that both straw-man solutions add significant overhead compared to the ideal performance. Copying to contiguous memory incurs cost proportional to the number of past KV-tokens. Applying multiple rounds of PagedAttention, on the other hand, gives up parallelization opportunities on prompt tokens, resulting in execution time linear to the number of tokens in the prompt. Pensieve’s kernel matches the ideal baseline. In fact, it has slightly better performance because we offload auxiliary data computing (like calculating the cumulative sum for the sequence length of a batch) to the CPU. Since each transformer layer in the model shares the same caching plan, these auxiliary data can be reused by all layers.
6.5. Effect of Unified Scheduling
Figure 13. Performance of serving Llama 2-13B with and without unified scheduling on ShareGPT dataset.
We evaluate whether using a unified scheduler for both prefill and generation phases is beneficial for performance. Figure 13 shows the performance of Pensieve with and without unified scheduling for Llama 2-13B on the ShareGPT dataset. Compared to processing prefill and generation phases separately, unifying them into a single execution step avoids having to execute the prefill phase with a small number of requests. As a result, Pensieve with unified scheduling achieves better throughput and latency.
6.6. Effect of the Eviction Policy
Figure 14. Performance of serving OPT-13B with different eviction policies on ShareGPT dataset.
We compare Pensieve’s caching policy against the classic LRU policy. We use OPT-13B as the workload for this evaluation. From Figure 14 shows that both policies exhibit similar performance until the workload approaches 3 requests per second, beyond which Pensieve’s eviction policy outperforms LRU. After analyzing the execution traces, we find that both policies have less than 80% cache hit rate, however, Pensieve’s policy has up to 4.4 percentage points higher CPU cache hit rate than LRU. On average, Pensieve’s policy reduces the number of recomputed KV-tokens by up to 14.6%.
6.7. Impact of User Think Time
Figure 15. Performance of serving Llama 2-13B with different average user reaction time on ShareGPT dataset.
Our experiments so far use the average user think time of 60 seconds. Figure 15 evaluates the impact of different average user think times on the performance of Pensieve using Llama 2-13B. Additionally, we also show vLLM with 600-second think time as a comparison point. As seen in Figure 15, the throughput of Pensieve decreases as the average user think time increases, causing past KV-tokens to drop from the cache at a higher rate. Even as user think time increases to 600 seconds, Pensieve still achieves better latency and throughput compared to vLLM, although the performance gap becomes smaller than that for smaller think times.
@@ 7. Related Works
LLM inference and serving systems.
Many systems have been recently developed to serve language models with better performance including vLLM (Kwon et al., 2023), Orca (Yu et al., 2022), TensorRT-LLM (NVIDIA, 2023), FasterTransformer (NVIDIA, 2021), LightSeq (Wang et al., 2021), DeepSpeed (Aminabadi et al., 2022) and  (Pope et al., 2023). These systems investigate different performance improvement opportunities than Pensieve. The wide range of proposed techniques include incremental decoding (NVIDIA, 2021), iteration-level batching (Yu et al., 2022; Gao et al., 2018), supporting paged KV cache over non-contiguous GPU memory (Kwon et al., 2023; NVIDIA, 2023), finding an efficient multi-device partitioning plan (Pope et al., 2023), better load balancing between prefill and generation to reduce pipeline parallelism bubbles (Agrawal et al., 2023; Holmes et al., 2024), kernel fusion (Dao et al., 2022, 2023; Rabe and Staats, 2021), speculative decoding (Chen et al., 2023; Leviathan et al., 2023), and quantization (Nagel et al., 2021; Gholami et al., 2022; Dettmers and Zettlemoyer, 2023; Xiao et al., 2023).
ServerlessLLM (Fu et al., 2024) proposes techniques to do fast model loading and live migration in order to serve requests in a serverless setup. DistServe (Zhong et al., 2024) advocates performing the prefill and generation phase of the same request on separate GPUs to better satisfy SLA. By contrast, Pensieve performs unified batching of both phases because we aim to optimize throughput instead of latency SLA.
The system of  (Pope et al., 2023) can keep a fixed-sized KV cache for each conversation in TPU device memory. Such caching is too rigid. By contrast, Pensieve manages the cache across all conversations and can also swap to CPU to relieve GPU memory pressure. Another recent system, CacheGen (Liu et al., 2023), addresses the problem of reducing the context-loading delay for long contexts stored in remote network storage. CacheGen proposes an adaptive compression scheme to accelerate the transfer of the KV cache over a bandwidth-limited and bursty network. As Pensieve only keeps the KV cache in local GPU and CPU memory, CacheGen’s techniques are orthogonal to those of Pensieve. For conversational chatbots, Pensieve using local cache can already bring significant performance gains.
Non-LLM specific DNN serving systems.
@@
Systems like TensorFlow Serving (Olston et al., 2017), Clipper (Crankshaw et al., 2017), NVIDIA Triton Inference Server (NVIDIA, 2018), Nexus (Shen et al., 2019), and InferLine (Crankshaw et al., 2020) serve as scheduling components of a general-purpose DNN serving system. They are mostly model-agnostic and execution backend-agnostic and apply general system techniques like batching, caching, and software pipelining to serve DNN models. Some are also in charge of properly provisioning compute resources to improve overall cluster efficiency. A few existing works target model-less serving and provide inference as a service: they automatically select models to meet the accuracy and latency requirements of a given user task. For example, INFaaS (Romero et al., 2021) and Tabi (Wang et al., 2023) serve requests with a small model and only re-route to a larger model when the output confidence score is low.
Techniques addressing the GPU memory limit.
These include GPU-CPU swapping, recomputation, and unified memory. Most of the systems described below are not specifically targeted for LLM serving.
GPU-CPU Swapping: SwapAdvisor (Huang et al., 2020) swaps weight and activation tensors for DNN training. It uses the dataflow graph to determine an optimal plan that involves operator execution order, memory allocation, and swapping. Zero-Offload (Ren et al., 2021) offloads optimizer state and gradients to the CPU during LLM training with a single GPU. DeepSpeed-ZeroInference (Aminabadi et al., 2022) and FlexGen (Sheng et al., 2023) use offloading to serve LLM with a weak GPU. DeepSpeed-ZeroInference offloads entire model weights to CPU or NVMe memory. FlexGen offloads currently unused model weights, activation, or KV cache to CPU memory or disk. As a result of frequent data movement and high disk latency, these two systems require a large batch size to hide the offloading latency. Therefore, they mainly target latency-insensitive applications. Neither system persists the KV cache across requests.
Recomputation: For DNN training, recomputing activation on the backward pass (Chen et al., 2016) is a popular technique used to reduce memory footprint. The decision for which tensors to compute is done at the granularity of layers (Chen et al., 2016), a group of operators within a layer (aka modules) (Korthikanti et al., 2023), or individual operators (Jain et al., 2020). Unlike these works, Pensieve’s fine-grained recomputation is done at the token(chunk)-level, to exploit the insight that earlier tokens in a sequence incur less recomputation cost due to the causal nature of attention.
Unified Memory and Direct Host Access: Swapping between CPU and GPU memory can also be achieved transparently through Unified Memory (NVIDIA, 2017a), which automatically triggers memory page migration between CPU and GPU using a page fault mechanism. The Direct Host Access feature allows a GPU kernel to directly read from CPU memory, thereby allowing a subset of the threads to execute as soon as their data is ready without waiting for the entire transfer to finish. It has been used to enable better overlap of data transfer and kernel computation when swapping in a model from CPU (Min et al., 2021) or accessing large graph neural network features on the CPU (Jeong et al., 2023). For our implementation of Pensieve, we choose not to use Unified Memory nor Direct Host Access because these mechanisms trigger memory transfer only when the data is accessed by a GPU kernel and we want to manage data movement explicitly to prefetch the KV cache. Furthermore, for Direct Host Access, since the copied data is not explicitly stored in GPU memory, it has to be loaded from the CPU again if it is repeatedly used. In our workload, the context cache is not only used to prefill the prompt tokens but also in every generation step, which makes Direct Host Access significantly less efficient.